Trains A and B are traveling in the same direction. Train A is traveling at 100 miles per hour and train B is traveling at 125 miles per hour. Train A passes a station at 11:10pm,if Train B passes the same train station at 11:22pm. At what time will train B catch up with train A?
Answer by htmentor(1343)

Train A has traveled for 12 minutes at 100 mph by the time train B reaches the station.
First convert 12 minutes to hours:
12 min * 1 hr/60 min = 1/5 hr
When train B reaches the station, the trains are separated by a distance
d = 1/5*100 = 20 miles
The relative speed of the two trains is 125 - 100 = 25 mph
So the time required for train B to catch train A is:
20 mi/25 mi/hr = 4/5 hr = 48 minutes
So the time will be 48 minutes after 11:22 pm, which is 12:10 am